OpenStudy (anonymous):

the flower garden has the shape of triangle 34ft of a perrenial border forms the hypotenuse of the triangle and one leg is 14 ft longer than the other leg. find the length of the legs

OpenStudy (anonymous):

The setup for this problem is a triangle that is (x+14)-(x)-34. I am assuming this is a right triangle since no angles are given. \[(x+14)^{2}+x^{2}=34^{2}\] Expanded: \[x^2+28x+196+x^2=1156\] Simplified: \[2x^2+28x+196=1156\] Set to 0: \[2x^2+28x-960=0\] Divide by 2 to make factoring easier: \[x^2+14x-480=0\] Factor: \[(x+30)(x-16)=0\] Therefore: \[x=-30,16\] Discard the negative value as side cannot be negative. Therefore the sides are: 30-16-34 Proof: \[30^2+16^2=34^2\] \[900+256=1156\]